OpenStudy (anonymous):

What is an equivalent form of 3 ----- 4-2i

OpenStudy (danjs):

I think you need to multiply top and bottom by 4+2i

OpenStudy (anonymous):

oh ok

OpenStudy (danjs):

u have it from there?

OpenStudy (anonymous):

doesn't it cancel?

OpenStudy (nincompoop):

it doesn't

OpenStudy (danjs):

use i^2 = -1

OpenStudy (anonymous):

oohh ok

OpenStudy (nincompoop):

multiplying by the conjugate doesn't necessarily cancel unless there is something to cancel.

OpenStudy (anonymous):

hmmm

OpenStudy (danjs):

3/(4-2i) = (12 + 6i)/ (16 +4i^2)

OpenStudy (danjs):

i^2=-1

OpenStudy (anonymous):

i see, so i becomes -1 regardless. and the answer is that equation?

OpenStudy (danjs):

put in -1 for i^2 and simplify the denominator

OpenStudy (anonymous):

because it doesn't tell me to solve it

OpenStudy (anonymous):

ok i get it

OpenStudy (danjs):

(12+6i)/)(16-4)

OpenStudy (anonymous):

-18/12

OpenStudy (anonymous):

?

OpenStudy (danjs):

It is the same as though you are using normal variables, just recall that i^2=-1 in your final answer

OpenStudy (anonymous):

Ok thank you!!

OpenStudy (danjs):

no the numorator has just i in it, you have to leave that alone

OpenStudy (danjs):

(12+6i)/12

OpenStudy (danjs):

now you can factor a 6 out of top and bottom

OpenStudy (danjs):

[6(2+i)] / (6*2) cancel the 6's

OpenStudy (danjs):

(2+i) / 2

OpenStudy (anonymous):

wow, you are awesome thank you

OpenStudy (danjs):

np, anytime